OpenStudy (anonymous):

the volume, in cubic centimeters, of a sphere whose diameter is 6 cm is?

OpenStudy (anonymous):

4/3pi*r^3=4/3pi*3^3=36pi cubic centimeters

OpenStudy (saifoo.khan):

use the volume formula.

OpenStudy (anonymous):

volume of a sphere is\[(4/3)(\pi)r^3\] so take half the diameter to find the radius, and substitute it for r in the formula

OpenStudy (anonymous):

i dont understand how u got 36 though

OpenStudy (anonymous):

(4/3)pi(3^3) =4*9pi =36pi

OpenStudy (anonymous):

so u rounded the 4?